what is the amount of power applied by the wind glass to a 20kg anchor if it is lifted 2-m high within 5s?
1
Expert's answer
2021-09-06T11:33:19-0400
"P=\dfrac{W}{t}=\dfrac{mgh}{t},""P=\dfrac{20\ kg\cdot10\ \dfrac{m}{s^2}\cdot2\ m}{5\ s}=80\ W."